OpenStudy (anonymous):

convert r=k/(4cos theta + 6sin theta) to cartesian

OpenStudy (anonymous):

Using T for theta x = r.cosT y = r.sinT you give r = k/(4cosT + 6 sinT) the denominator become: 4x/r + 6y/r = (4x+6y)/r so r = k divided by (4x+6y)/r to divide by a fraction, you flip it & multiply so r = k times r/(4x+6y) r = kr/(4x+6y) cancel the common factor of r 1 = k/(4x+6y) giving 4x+6y = k